OpenStudy (anonymous):

Eliminate the parameter. x = square root of t, y = 4t + 1

OpenStudy (agent0smith):

\[\large x = \sqrt t\]just replace t in the y = 4t + 1 with x^2

OpenStudy (agent0smith):

since \[\large x = \sqrt t \text { then } x^2 = t\]

OpenStudy (anonymous):

parameter is t, then you have to eliminate t, so just replace \[t = x^2\] into y = 4t+1, like @agent0smith said above

OpenStudy (anonymous):

so it would be y=4x^2+1

OpenStudy (agent0smith):

Yep.